The work function (Φ) for a metal is 6.10×10-19 J. What is the longest wavelength of electromagnetic radiation that can eject an electron from the surface of a piece of the metal? Answer in nm.
1
Expert's answer
2019-10-01T06:03:05-0400

Answer: 326 nm


Solution:


"E=\frac{hc}{\lambda}"


"\lambda=\frac{hc}E"


"\lambda=\frac{6.626×10^{-34}Js\;×\;3×10^8 m/s}{6.10×10^{-19}J}=3.26×10^{-6}m=326\,nm"